OpenStudy (anonymous):

Integrate f'(g (5x)) g'(5x)

OpenStudy (solomonzelman):

more info, or is that it?

jimthompson5910 (jim_thompson5910):

Let h(x) = f(g(5x))

jimthompson5910 (jim_thompson5910):

if we compute h ' (x), we get h(x) = f(g(5x)) h ' (x) = f ' (g(5x)) * g ' (5x) * 5 ... use chain rule twice h ' (x) = 5f ' (g(5x)) * g ' (5x) so h ' (x) is very close to what we're given, which is f'(g (5x)) g'(5x). The only difference really is that 5 out front for h ' (x).

jimthompson5910 (jim_thompson5910):

\[\Large h^{\prime}(x) = 5f^{\prime}(g(5x))*g^{\prime}(5x)\] \[\Large \frac{1}{5}h^{\prime}(x) = f^{\prime}(g(5x))*g^{\prime}(5x)\] \[\Large \int\frac{1}{5}h^{\prime}(x)dx = \int f^{\prime}(g(5x))*g^{\prime}(5x)dx\] I'm sure you see what to do from here.